OpenStudy (anonymous):

What is the equation of the line of symmetry for the graph of y = 3(x + 2)^2 - 4? x = -4 x = -2 x = 2 x = 3

OpenStudy (anonymous):

here we are looking for the x coordinate of the vertex. y=a(x -h)^2 +k where the vertex is given by (h,k) so with your equation what will the x symmetry line be?

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

close if I put x=2 back into the equation I will get this y=a(x -2)^2 +k but we have y = 3(x + 2)^2 - 4 so x=-2. to check put it back into the equation y=a(x -(-2))^2 +k y=a(x +2)^2 +k

OpenStudy (anonymous):

ohhh i see... thank you

OpenStudy (anonymous):

no worries :)

OpenStudy (anonymous):

also. the vertex = -b/2a, that will give you the line of symmetry x coord also since the vertex is directly in the middle :)